OpenStudy (anonymous):

Choose the correct simplification of the expression g^5h^4 over g^2h^3

OpenStudy (anonymous):

\[\frac{ g^5h^4}{g^2h^3 } = g^{5-2} h^{4-3} =g^3h\]